OpenStudy (anonymous):

What is the heating capacity, in BTUs, of a cylindrical storage tank of natural gas with a radius of 3 feet and a height of 10 feet? Assume that 1 cubic foot of propane contains 2,516 BTUs and 1 cubic foot of natural gas contains 1,030 BTUs. Use 3.14 for pi, and round your answer to the nearest BTU.

OpenStudy (anonymous):

volume of tank is given by = 3.14 *3*3*10 cubic foot So 3.14 *3*3*10 cubic foot of natural gas = 3.14 *3*3*10 *1,030 BTUs